### v1

# Enhanced quantum radiation with flying-focus laser pulses

Abstract
The emission of a photon by an electron in an intense laser field is one of the most fundamental processes in electrodynamics and underlies the many applications that utilize high-energy photon beams. This process is typically studied for electrons colliding head-on with a stationary-focus laser pulse. Here, we show that the energy lost by electrons and the yield of emitted photons can be substantially increased by replacing a stationary-focus pulse with an equal-energy flying-focus pulse whose focus co-propagates with the electrons. These advantages of the flying focus are a result of operating in the quantum regime of the interaction, where the energy loss and photon yield scale more favorably with the interaction time than the laser intensity. Simulations of 10 GeV electrons colliding with 10 J pulses demonstrate these advantages and predict a 5×5$\times$5 × increase in the yield of 1-20 MeV photons with a flying focus pulse, which would impact applications in medicine, material science, and nuclear physics.
Modern high-power lasers enable direct probing of strong-field quantum electrodymanics with optical light [1, 2, 3, 4, 5, 6]. Strong-field quantum electrodynamics (SFQED) refers to processes that occur in electromagnetic fields with amplitudes of the order of the Schwinger critical field $F_{\text{cr}}=m^{2}/|e|=1.3\times 10^{16}$ V/cm = $4.4\times 10^{13}$ G [7, 8, 9, 10]. Here, $e<0$ is the electron charge, $m$ its mass, and units $\hbar=c=\epsilon_{0}=1$ are employed throughout. The corresponding intensity is $I_{\text{cr}}=4.6\times 10^{29}$ W/cm2. To date, the highest intensity achieved by a high-power laser in the lab frame is about six orders of magnitude below this value [11]. Nevertheless, the peak intensity experienced by an ultrarelativistic electron in its rest frame can readily exceed $I_{\text{cr}}$. The key parameter that determines the strength of a SFQED process is $\chi=\sqrt{|F^{\mu\nu}p_{\nu}|^{2}}/mF_{\text{cr}}$, where $p_{\nu}$ is the electron four-momentum and $F^{\mu\nu}$ is the electromagnetic tensor of the laser field. Because $\chi$ is a Lorentz-invariant quantity, it is directly proportional to the electric field experienced by an electron in its rest frame. As a result, a value of $\chi\gtrsim 1$ can be achieved experimentally by colliding available multi-GeV electron beams [12, 13, 14] with existing high-power laser pulses.
Space-time structured laser pulses, such as the flying focus (FF), offer a new paradigm for probing SFQED processes. Flying focus pulses feature a programmable velocity focus that moves independently of the group velocity [15, 16, 17]. The moving intensity peak formed by the flying focus can travel distances much longer than a Rayleigh range while maintaining a near-constant spatiotemporal profile [16, 18, 17, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Experimental configurations and conceptual proposals for producing FF pulses have used chromatic optics [16, 17, 23], axiparabola-echelon pairs [22, 30, 31, 26], and nonlinear optical processes [20, 24]. In the context of high-field physics, a FF pulse with an intensity peak that moves at the speed of light in the opposite direction of its phase fronts allows for extended interaction times with high energy particles that collide head-on with the phase fronts. This configuration was first suggested to accentuate signatures of the transverse-formation-length in the emission of radiation [32]. The same configuration was later proposed in Refs. [33, 34] to facilitate the detection of classical radiation-reaction and quantum vacuum-birefringence. The observed effect in both cases, i.e., the classical energy loss of the electron beam and polarization rotation of an x-ray probe beam, scaled with the energy of the laser pulse [7, 8, 9, 10]. Thus, the extended interaction time afforded by the FF enabled the use of much lower and more controllable powers than ultrashort, stationary-focus Gaussian (SFG) pulses. The magnitude of the effect produced by the FF and SFG was, however, the same.
In this letter, we show that a FF pulse can enhance observable effects of quantum radiation reaction when compared to a SFG with the same energy. More specifically, FF pulses result in a greater loss of electron energy and a larger yield of photons in the energy range $1\text{-}20\;\text{MeV}$ than SFG pulses. This is because in the quantum regime ($\chi{\sim}1$) both the electron energy loss and photon yield scale more favorably with the interaction time than the laser intensity. A larger yield of photons in the $1\text{-}20\;\text{MeV}$ would impact several applications, including radiosurgery [35, 36], photo-transmutation of long-lived nuclear waste and production of medical isotopes [37, 38, 39], and investigation of the structure of materials by nuclear resonance fluorescence [40, 41]. While many methods have been proposed to increase the yield of high-energy photons (see e.g., [42, 43, 44, 45, 46, 47]), here, we show something qualitatively different: how space-time structured light can significantly enhance the emission probability of the underlying process.
The basic SFQED process describing radiation by an ultrarelativistic electron beam in the field of a laser pulse is nonlinear Compton scattering. Here, this process will be investigated within the locally-constant field approximation (LCFA) [7, 8, 9, 10]. The LCFA is a useful approximation to study SFQED phenomena because it allows one to compute the probability of a process in an arbitrary electromagnetic field using a known expression for the probability in a constant crossed field. A constant crossed field is a space and time-independent electromagnetic field $(\bm{E},\bm{B})$ that satisfies $|\bm{E}|=|\bm{B}|$ and $\bm{E}\cdot\bm{B}=0$ [48, 49, 50]. The LCFA is valid provided that: (a) ultrarelativistic charges move fast enough to experience an arbitrary electromagnetic field as a crossed field in their rest frame [51] and (b) the formation length of a process like nonlinear Compton scattering is much smaller than the wavelength of the laser pulse $\lambda_{0}=2\pi/\omega_{0}$, where $\omega_{0}$ is central frequency of the pulse. This latter condition is typically valid when the dimensionless field strength $\xi_{0}=|e|E_{0}/m\omega_{0}\gg 1$, where $E_{0}$ is the electric field amplitude of the pulse [48, 50].
Within the LCFA, the average energy $\mathcal{E}_{\gamma}$ radiated by an electron per unit time in a background electromagnetic field via the emission of a single photon can be calculated using the Baier-Katkov interpolation formula [50]
$\frac{d\mathcal{E}_{\gamma}}{dt}\approx\frac{2\alpha m^{2}\chi^{2}/3}{[1+4.8(1+\chi)\ln(1+1.7\chi)+2.44\chi^{2}]^{2/3}},$ (1)
where $\alpha\approx 1/137$ is the fine-structure constant and $\chi$ is evaluated using the local values of $p_{\nu}$ and $F^{\mu\nu}$. The interpolation has an accuracy better than 2% for all values of $\chi$ [50]. When the successive emission of each photon is uncorrelated, $-d\mathcal{E}_{\gamma}/dt$ is equal to the average energy lost by an electron per unit time, i.e., $d\mathcal{E}/dt=-d\mathcal{E}_{\gamma}/dt$, where $\mathcal{E}=\mathcal{E}(t)$ is the electron energy at time $t$ [52, 53]. Thus, Eq. (1) can be used to calculate the relative electron energy loss $\zeta$ defined as
$\zeta\equiv\frac{\mathcal{E}_{0}-\mathcal{E}_{F}}{\mathcal{E}_{0}},$ (2)
where $\mathcal{E}_{0}$ is the initial electron energy and $\mathcal{E}_{F}$ is its energy after the interaction with the laser pulse. While correlated photon emission can be important in the quantum regime [53], it is negligible for the parameters considered here. This is verified in Section I of the Supplemental Material (SM) by comparing the results of Eq. (1) to a full quantum mechanical approach (see also Ref. [54]).
Figure 1: Relative energy loss $\zeta$ as a function of the interaction time for an electron colliding head-on and on-axis with equal-energy FF and SFG pulses. From top to bottom, the lines correspond to initial electron energies of 10, 5, 2, 1, and 0.5 GeV . Both the FF and SFG pulses had a $\lambda_{0}=1\ \mu$m wavelength, $\sigma_{0}=1.5\ \mu$m focused spot size, maximum quantum nonlinearity parameter $\chi_{0}\approx 0.1\xi_{0}\mathcal{E}_{0}$(10 GeV), and energies of either (a) $U=1\;\text{J}$ or (b) $U=10\;\text{J}$. For the SFG, $z_{R}=42$ fs. To the right of the dash-dotted vertical line, the results may have errors larger than 4% due to the breakdown of the LCFA. Within the quantum regime $\chi_{0}\sim 1$ and LCFA, the maximum energy loss achieved by a FF pulse is substantially greater than the maximum achieved by a SFG pulse, especially for $U=1\;\text{J}$.
The advantage of using a FF pulse as compared to a SFG pulse with the same energy will be demonstrated by numerically integrating Eq. (1). An analytical treatment is also possible by making a few simplifying assumptions: (1) $\mathcal{E}\gg m\xi_{0}$, which guarantees that, for an electron initially counterpropagating with respect to the laser pulse, the transverse electron motion and pulse structure can be ignored; (2) The focal point of the laser pulse is located at $x=y=0$ and the electron moves along the negative $z$ axis at an ultrarelativistic velocity such that $z(t)\approx z(0)-t$; (3) The temporal profile of the pulse has a square shape equal to unity with a duration $2t_{\text{int}}$, which intersects the electron from $z=t_{\text{int}}/2$ to $z=-t_{\text{int}}/2$.
In the classical regime ($\chi\ll 1$), Eq. (1) can be approximated as $d\mathcal{E}/dt\approx-2\alpha m^{2}\chi^{2}/3$. When interacting with the FF pulse, the electron experiences a field with a near-constant amplitude $\xi_{0}$ such that $\chi_{\text{FF}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}$. The relative energy loss [33] is then
$\zeta_{\text{C,FF}}=\frac{\kappa_{\text{C,FF}}}{1+\kappa_{\text{C,FF}}},\>\>\kappa_{\text{C,FF}}\approx 2\frac{U(\text{J})\mathcal{E}_{0}(\text{GeV})}{[\sigma_{0}(\mu\text{m})]^{2}}\,,$ (3)
where $U=\pi E_{0}^{2}\sigma_{0}^{2}t_{\text{int}}/2$ is the pulse energy [55] and $\sigma_{0}$ its focused spot size. This result shows that the energy loss in the classical regime depends on the fluence of the laser pulse, i.e., it is $\propto U/\sigma_{0}^{2}$, but is independent of the interaction time $t_{\text{int}}$ [56]. For a SFG pulse, the finite Rayleigh range $z_{R}=\omega_{0}\sigma_{0}^{2}/2$ can be taken into account by substituting $\xi_{0}\rightarrow\xi_{0}/\sqrt{1+z^{2}/z_{R}^{2}}=\xi_{0}/\sqrt{1+t^{2}/z_{R}^{2}}$ in the expression for $\chi$, such that $\chi_{\text{SFG}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}\sqrt{1+t^{2}/z_{R}^{2}}$. As a result, $\kappa_{\text{C,SFG}}=\kappa_{\text{C,FF}}\arctan(\rho)/\rho$, with $\rho\equiv t_{\text{int}}/2z_{R}$ [33]. Thus, for a fixed laser energy, a FF pulse and a short SFG pulse with $\rho\ll 1$ induce the same energy loss in the classical regime.
In the quantum regime ($\chi\gtrsim 1$), the emission rate [Eq. (1)] has a weaker dependence on the field amplitude than in the classical regime. This suggests that, in the quantum regime, a long FF pulse can induce more energy loss than a short SFG pulse with the same energy. This reasoning can be verified analytically for FF pulses in the limit $\chi\gg 1$, where $d\mathcal{E}/dt\propto-\chi^{2/3}$ and
$\mathcal{E}_{0}^{1/3}-\mathcal{E}_{F}^{1/3}\propto U^{1/3}t_{\text{int}}^{2/3}.$ (4)
Thus, for a fixed pulse energy $U$, higher energy losses can be achieved by extending the interaction time $t_{\text{int}}$.
Figure 1 confirms the above expectations by showing the relative energy loss as a function of the interaction time for electrons colliding head-on with either FF pulses (blue solid lines) or SFG pulses (dashed orange lines) with energies of (a) 1 J and (b) 10 J. The solid and dashed lines were obtained by numerically integrating Eq. (1), with $\chi$ equal to $\chi_{\text{FF}}$ or $\chi_{\text{SFG}}$, respectively. In both cases, the maximum value of the quantum nonlinearity parameter is $\chi_{0}=2\mathcal{E}_{0}\omega_{0}\xi_{0}/m^{2}\approx 0.1\,\xi_{0}\mathcal{E}_{0}(10\;\text{GeV})$. For $U=$ 1 J, the maximum energy loss achieved by a FF pulse is approximately $30\%$ higher than that of an SFG pulse regardless of the initial electron energy. For $U=$ 10 J and $\mathcal{E}_{0}=$ 0.5 GeV, the maximum energy loss achieved by a FF pulse is $12\%$ higher. In this case and for larger electron energies, the enhanced energy loss in the FF pulses is reduced because the electron loses almost all of its energy whether it interacts with a FF or SFG pulse (see also the inset of Fig. 2a discussed below).
Figure 1 also shows that the energy loss in SFG pulses is suppressed for either very short pulses ($\rho\ll 1$) or very long pulses ($\rho\gg 1$). Shorter duration SFG pulses have higher peak intensities, but, as discussed above, the emission rate has a weaker scaling with intensity than it does with duration in the quantum regime [57]. With long duration SFG pulses, the electron spends a large fraction of the interaction time outside of the confocal region where the intensity is lower. This contrasts the FF pulses where the electron spends the entire interaction near the focus where the intensity is high. As a result, the energy loss increases monotonically with the interaction time until it approaches the classical limit [green dotted lines evaluated using Eq. (3)]: for fixed pulse energy, $\chi_{0}\propto t_{\text{int}}^{-1/2}$. The energy loss in the FF and SFG pulses coincide at short interaction times because the two pulses have the same amplitude and the finite Rayleigh range of the SFG pulse does not play a role ($\rho\ll 1$).
The results presented in Fig. 1 were verified by evolving electrons in the full FF and SFG fields with a Monte Carlo radiation emission that takes into account the correlations in multiple photon emission (see Section I of SM [58]). Notwithstanding, some remarks are in order to address the validity of the LCFA. The LCFA overestimates the exact QED emission spectrum for sufficiently low-energy photons [59, 60, 61]. The key parameter is [59]
$\eta_{\text{LCFA}}=\frac{\mathcal{E}-\omega}{\omega}\frac{\chi}{\xi^{3}}\,,$ (5)
where $\omega$ is the energy of the emitted photon and $\xi$ and $\chi$ are the classical and quantum nonlinearity parameters at the instant of emission. The condition $\eta_{\text{LCFA}}^{2/3}\ll 1$ ensures that corrections to the LCFA are small [60]. Equation (1) is derived by integrating the emitted photon spectrum over the photon energy. For $\chi\ll 1$, the largest contribution to the integral comes from energies $\omega\lesssim\chi\mathcal{E}\ll\mathcal{E}$; for $\chi\sim 1$, the largest contribution comes from energies $\omega\lesssim\mathcal{E}$. As a result, the validity of the LCFA is determined by the low-energy photons and the condition $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi^{2}\ll 1$. The vertical line in Fig. 1 corresponds to the choice $\xi_{0}=5$ [62] so that $\eta_{\text{LCFA}}^{2/3}=1/\xi_{0}^{2}=0.04$. Note that the expression for the classical energy loss does not rely on the validity of the LCFA, which explains why in the FF case the curves tend to the classical result for longer and longer interaction times ($\chi_{0}\propto t_{\text{int}}^{-1/2}$ for fixed energy).
Figure 2: Relative energy loss $\zeta$ as a function of the initial electron energy for electrons colliding head-on and on-axis with equal-energy FF and SFG pulses. In (a), the pulses have an energy $U=$ 1 J, 10 J, or 100 J and a spot size $\sigma_{0}=$1.5 $\mu$m. In (b), $U=$ 10 J and the spot size is $\sigma_{0}=$ 1.5 $\mu$m or 4 $\mu$m. The color scale shows the maximum quantum nonlinearity parameter. Above $\mathcal{E}\approx$ 3 GeV, both the FF and SFG are in the quantum regime. The inset shows the relative improvement in the energy loss afforded by the FF pulses. At $U=$ 1 J, the FF results in ${\sim}30\%$ more loss than the SFG. At $U=$ 10 J, both the FF and SFG begin to approach $100\%$ loss, which limits the relative improvement achievable with a FF pulse.
To further illustrate the advantage of FF pulses in the quantum regime of radiation reaction, Fig. 2 compares the maximum energy loss induced by FF pulses (the longest $t_{\mathrm{int}}$ satisfying $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi_{0}^{2}<0.04$) to that of SFG pulses (the maxima of the orange dashed curves in Fig. 1). For almost all initial electron energies $\mathcal{E}_{0}$, the FF pulses (solid lines) result in a larger loss of electron energy than the SFG pulses (dashed lines). However, as the pulse energy increases, the electron energy loss for both the FF and SFG pulses approaches unity (Fig. 2a), which limits the relative advantage of the FF pulses (Fig. 2a inset). In addition, both pulses achieve higher energy losses with smaller spot sizes (Fig. 2b) [see also Eq. (3) for the classical regime]. Thus, defocusing an SFG pulse to increase the interaction time would not improve its results.
The saturation of the electron energy loss at higher pulse energies suggests that an observable without a theoretical upper limit would better elucidate the benefits of FF pulses in these conditions. One such observable, that is also important for applications, is the probability $P$ of photon emission per unit time. Multiplying the probability by the number of electrons colliding with the laser pulse provides the average number of emitted photons. This number is exact for uncorrelated photon emission [52, 53] and is an excellent approximation when the contribution from correlated photon emission is negligible, as is the case here (see Section I of the SM [58]).
The interpolation formula for the probability of single photon emission per unit time within the LCFA is given by [50]
$\frac{dP}{dt}\approx\frac{1.45\alpha m^{2}\chi}{\mathcal{E}(1+7.2\chi+\chi^{2})^{1/6}}\,,$ (6)
which is better than 2% accurate for all values of $\chi$. In the classical regime ($\chi\ll 1$), $dP/dt\propto\chi/\mathcal{E}$; while in the quantum regime ($\chi\gg 1$), $dP/dt\propto\chi^{2/3}/\mathcal{E}$. These proportionalities indicate that the emission probability $P$ scales more favorably with the interaction time than the laser intensity in both regimes. Thus, for fixed collision parameters (i.e., pulse energy, electron energy, focusing), FF pulses will have an advantage over SFG pulses regardless of $\chi$.
For definiteness, consider the most favorable conditions for a SFG pulse, which is the classical regime where $\chi\ll 1$. Maximizing the emission probability for a SFG pulse requires matching the interaction time $t_{\text{int}}$ to the Rayleigh range $z_{R}$, so that electrons spend much of the interaction near the focus where the intensity is high. The optimum is $t_{\text{int}}\approx 6z_{R}$, which yields $P_{\text{C,SFG}}\propto U^{1/2}$ (see Section II of SM [58] for derivation). With a FF pulse, the focus moves with the electrons, so the entire interaction time is spent near the focus where the intensity is high. As a result, the interaction time can be independent of the FF Rayleigh range $z_{R,\text{FF}}\equiv\omega_{0}\sigma_{0}^{2}$. In this case, $P_{\text{C,FF}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}U^{1/2}$, and the optimum is achieved by making the interaction time as long as possible. The ratio $P_{\text{C,\text{FF}}}/P_{\text{C,SFG}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}$ reveals that, for fixed laser pulse energy, a FF pulse will have an advantage over a SFG pulse when the interaction time of the FF is longer than its Rayleigh range.
The advantage of FF pulses is confirmed in Fig. 3, which compares the number of photons $\hat{N}_{\gamma}$ emitted per unit photon energy per electron from the collision of $\mathcal{E}_{0}=$ 10 GeV electrons with $U=$ 10 J pulses. The parameters were chosen to be experimentally feasible [6]: the wavelength and spot size of the laser pulses were $\lambda_{0}=1\ \mu$m and $\sigma_{0}=2\ \mu$m, respectively, and the electron beam was initialized with normally distributed transverse positions with a variance of $0.75\ \mu$m, a divergence of 1 mrad, and a length of $5\lambda_{0}$. The remaining parameters were determined by maximizing the photon yields from the SFG and FF pulses while still satisfying the LCFA: The FF pulse (solid line) had $\xi_{0}=6$ and $t_{\text{int}}=1.61$ ps, and the SFG pulse (dashed line) had $\xi_{0}=16$ and $t_{\text{int}}=227$ fs. Note that the choice of $\xi_{0}=6$ in the FF case (as opposed to $\xi_{0}=5$ used in Fig. 1) ensures the validity of the LCFA for off-axis electrons that experience lower field strengths. The photon spectra were generated using a Monte Carlo code (see Section I of SM [58]).
Figure 3 shows that the FF pulse results in more photons per electron than the SFG pulse by a factor of 7 in the 1 - 2 MeV range, a factor of 5 in the 2 - 10 MeV, and a factor of 4 in 10 - 20 MeV range. Based on these results, the collision of a 100 pC beam of $\mathcal{E}_{0}$ = 10 GeV electrons with a FF pulse could produce $5\times 10^{10}$ photons in the 1 - 20 MeV range, compared to only $9\times 10^{9}$ photons with a SFG pulse. Moreover, despite its much lower amplitude ($\xi_{0}=6$), the FF pulse still produces a substantial number of high-energy photons ($\omega\sim\mathcal{E}_{0}$) nearly equal to that produced by the higher amplitude ($\xi_{0}=16$) SFG pulse.
Figure 3: Spectra of emitted photons from the head-on collision of $\mathcal{E}_{0}=10$ GeV electrons with $U=10$ J laser pulses. The inset shows the number of photons emitted as a function of energy and $\eta_{\text{LCFA}}^{2/3}$. The number of photons emitted per electron in each shaded interval is displayed in the bottom left. The FF pulse (blue solid) results in approximately 5×5$\times$5 × more photons in the 1-20 MeV range than the SFG pulse (orange dashed).
Because the advantage of the FF pulse occurs in the low-energy part of the spectrum, it is important to verify that the breakdown of the LCFA at low photon energies does not invalidate the results [see Eq. (5) and Refs. [59, 60, 61]]. To assess possible deviations from the LCFA in this part of the spectrum, simulations were performed with the particle-in-cell (PIC) code Smilei [63] including the SFQED-tookit [64], which implements corrections beyond the LCFA [60]. These simulations are in excellent agreement with the Monte Carlo results presented in Fig. 3 (see Section III of SM [58]). As additional verification, the inset of Fig. 3 shows that almost all photons in the 1 - 20 MeV range are emitted with $\eta_{\text{LCFA}}^{2/3}<0.1$.
Finally, the Monte Carlo code used to generate the photon spectra in Fig. 3 does not account for the decay of the emitted photons into electron-positron pairs, which would also radiate [65]. As demonstrated in Section IV of the SM [58], pair production can be safely neglected for both the SFG and FF pulses considered in Fig. 3. At higher field strengths ($\xi_{0}\gtrsim 40$), pair production in a SFG pulse would deplete the most energetic photons and slightly increase the low-energy photon yield, but the overall energy loss and photon yield would be suppressed (see Section III of the SM [58]).
In conclusion, we have shown that a flying focus (FF) pulse can significantly enhance the signatures of quantum radiation reaction when compared to a stationary focus Gaussian (SFG) pulse with the same energy. First, an electron can radiate significantly more energy in a FF pulse than in a SFG pulse. Second, and perhaps more importantly, a FF pulse enhances the yield of emitted photons, especially in the low-energy part of the spectrum. As an example, a 10 GeV electron colliding with a 10 J flying focus pulse produces 5×5$\times$5 × more 1 - 20 MeV photons than with a SFG pulse. From a practical standpoint, this enhancement could benefit applications, such as radiosurgery, nuclear chemistry, and non-destructive imaging.
Acknowledgements.
The authors thank Samuele Montefiori and Matteo Tamburini for kindly sharing a development version of their SFQED-toolkit module [64] for SMILEI and providing support. This project has received funding from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101105246-STEFF. This material is based upon work supported by the U.S. Department of Energy [National Nuclear Security Administration] University of Rochester “National Inertial Confinement Fusion Program” under Award Number DE-NA0004144 and U.S. Department of Energy, Office of Science, under Award Number DE-SC0021057. This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.