### v2

# Enhanced quantum radiation with flying-focus laser pulses

Abstract
The emission of a photon by an electron in an intense laser field is one of the most fundamental processes in electrodynamics and underlies the many applications that utilize high-energy photon beams. This process is typically studied for electrons colliding head-on with a stationary-focus laser pulse. Here, we show that the energy lost by electrons in the quantum regime and the yield of emitted photons can be substantially increased by replacing a stationary-focus pulse with an equal-energy flying-focus pulse whose focus co-propagates with the electrons. These advantages of the flying focus result from the energy loss and the photon yield scaling more favorably with the interaction time than the laser intensity in the quantum regime, with the latter also holding in the classical regime. Monte Carlo simulations of electrons colliding with equal-energy stationary and flying-focus laser pulses demonstrate these advantages.
I Introduction
Modern high-power lasers enable direct probing of strong-field quantum electrodynamics with optical light [E144:1996enr, Burke:1997ew, Cole:2017zca, Poder:2017dpw, mirzaie2024, los2024observation]. Strong-field quantum electrodynamics (SFQED) refers to processes that occur in electromagnetic fields with amplitudes of the order of the Schwinger critical field $F_{\text{cr}}=m^{2}/|e|=1.3\times 10^{16}$ V/cm = $4.4\times 10^{13}$ G [DiPiazza:2011tq, Burton:2014wsa, Gonoskov:2021hwf, Fedotov:2022ely]. Here, $e<0$ is the electron charge, $m$ its mass, and units $\hbar=c=\epsilon_{0}=1$ are employed throughout. The corresponding intensity is $I_{\text{cr}}=4.6\times 10^{29}$ W/cm2. To date, the highest intensity achieved by a high-power laser in the laboratory frame is about six orders of magnitude below this value [Yoon:2021ony]. Nevertheless, the peak intensity experienced by an ultrarelativistic electron in its rest frame can readily exceed $I_{\text{cr}}$. The key parameter that determines the strength of a SFQED process is $\chi=\sqrt{-(F^{\mu\nu}p_{\nu})^{2}}/mF_{\text{cr}}$, where $p_{\nu}$ is the electron four-momentum, $F^{\mu\nu}$ is the electromagnetic tensor of the laser field, and the metric is $\eta^{\mu\nu}=\text{diag}(+1,-1,-1,-1)$. Because $\chi$ is a Lorentz-invariant quantity, it is directly proportional to the electric field experienced by an electron in its rest frame. As a result, a value of $\chi\gtrsim 1$ can be achieved experimentally by colliding available beams of multi-GeV electrons [Gonsalves:2019wnc, Abramowicz_2019, Meuren_2020, Picksley:2024cdd] with existing high-power laser pulses.
Space-time structured laser pulses, such as the flying focus (FF), offer a new paradigm for probing SFQED processes. FF pulses feature a programmable velocity focus that moves independently of the group velocity [Sainte-Marie_2017, Froula_2018, Turnbull_2018]. The moving intensity peak formed by the FF can travel distances much longer than a Rayleigh range while maintaining a near-constant spatiotemporal profile [Froula_2018, Palastro_2018, Turnbull_2018, Howard_2019, simpson2020nonlinear, Ramsey_2020, Palastro:2020gcl, Jolly:2020, simpson2022spatiotemporal, Ramsey_2022, pigeon2024ultrabroadband, Formanek:2023zia, ramsey2023exact, gong2024laser]. Experimental configurations and conceptual proposals for producing FF pulses have used chromatic optics [Froula_2018, Turnbull_2018, Jolly:2020], axiparabola-echelon pairs [Palastro:2020gcl, smartsev2019axiparabola, oubrerie2022axiparabola, pigeon2024ultrabroadband], and nonlinear optical processes [simpson2020nonlinear, simpson2022spatiotemporal]. In the context of high-field physics, a FF pulse with an intensity peak that moves at the speed of light in the opposite direction of its phase fronts allows for extended interaction times with high-energy particles that collide head-on with the phase fronts. This configuration was first suggested to accentuate signatures of the transverse-formation-length in the emission of radiation [DiPiazza:2020wxp]. The same configuration was later proposed in Refs. [Formanek:2021bpw, Formanek:2023mkx] to facilitate the detection of classical radiation-reaction and vacuum birefringence. The observed effect in both cases, i.e., the classical energy loss of the electron beam and polarization rotation of an x-ray probe beam, scaled with the energy of the laser pulse [DiPiazza:2011tq, Burton:2014wsa, Gonoskov:2021hwf, Fedotov:2022ely]. Thus, the extended interaction time afforded by the FF enabled the use of much lower and more controllable powers than ultrashort, stationary-focus Gaussian (SFG) pulses. The magnitude of the effect produced by the FF and SFG was, however, the same.
In this letter, we show that a FF pulse can significantly enhance observable effects of quantum radiation reaction when compared to a SFG with the same energy. More specifically, FF pulses result in a greater loss of electron energy and a larger yield of photons in the 1–20 MeV range than SFG pulses. This is because the electron energy loss scales more favorably with the interaction time than the laser intensity in the quantum regime ($\chi\gtrsim 1$), while the photon yield scales more favorably with the interaction time in both the classical and quantum regimes. A larger yield of photons in the 1–20 MeV range could impact several applications, including radiosurgery [girolami1996photon, weeks1997compton], photo-transmutation for treatment of long-lived nuclear waste and production of medical isotopes [wang2017transmutation, irani2014gamma, luo2016production], and investigations into the structure of materials using nuclear resonance fluorescence [kneissl1996investigation, albert2011design]. While many methods have been proposed to increase the yield of high-energy photons (see e.g., [powers2014quasi, vyskovcil2020inverse, luedtke2021creating, morris2021highly, hadjisolomou2023gamma, yu2024compact]), here, we show something qualitatively different: how space-time structured light can significantly enhance the emission probability of the underlying fundamental process.
II Radiation energy loss
The basic SFQED process describing radiation by an ultrarelativistic electron beam in the field of a laser pulse is nonlinear Compton scattering. Here, this process will be investigated within the locally-constant field approximation (LCFA) [DiPiazza:2011tq, Burton:2014wsa, Gonoskov:2021hwf, Fedotov:2022ely]. The LCFA is a useful approximation to study SFQED phenomena because it allows one to compute the probability of a process in an arbitrary electromagnetic field using a known expression for the probability in a constant crossed field. A constant crossed field is a space and time-independent electromagnetic field $(\bm{E},\bm{B})$ that satisfies $|\bm{E}|=|\bm{B}|$ and $\bm{E}\cdot\bm{B}=0$ [Ritus:1985vta, Reiss:1962nhe, Baier:1998vh]. The LCFA is valid provided that: (a) ultrarelativistic charges move fast enough to experience an arbitrary electromagnetic field as an approximate crossed field in their rest frame [Jackson_b_1975] and (b) the formation length of a process like nonlinear Compton scattering is much smaller than the wavelength of the laser pulse $\lambda_{0}=2\pi/\omega_{0}$, where $\omega_{0}$ is the central angular frequency of the pulse. This latter condition is typically satisfied when the dimensionless field strength $\xi_{0}=|e|E_{0}/m\omega_{0}\gg 1$, where $E_{0}$ is the electric field amplitude of the pulse [Ritus:1985vta, Baier:1998vh].
Within the LCFA, the average energy $\mathcal{E}_{\gamma}$ radiated by an electron per unit time in a background electromagnetic field through emission of a single photon can be calculated using the Baier-Katkov interpolation formula [Baier:1998vh]
$\frac{d\mathcal{E}_{\gamma}}{dt}\approx\frac{2\alpha m^{2}\chi^{2}/3}{[1+4.8(1+\chi)\ln(1+1.7\chi)+2.44\chi^{2}]^{2/3}},$ (1)
where $\alpha\approx 1/137$ is the fine-structure constant and $\chi$ is evaluated using the local values of $p_{\nu}$ and $F^{\mu\nu}$. The interpolation has an accuracy better than 2% for all values of $\chi$ [Baier:1998vh]. In the regime under consideration where the electron is ultrarelativistic ($\xi_{0}\gg 1$ and $\chi\sim 1$), the likelihood of multiple photon emission within a formation length is suppressed. Further, the work done by a laser pulse within a formation length is negligible compared to the energy lost to photon emission. Thus, the quantity $d\mathcal{E}_{\gamma}/dt$ is approximately equal to the average energy lost by an electron per unit time, i.e., $d\mathcal{E}/dt=-d\mathcal{E}_{\gamma}/dt$, where $\mathcal{E}=\mathcal{E}(t)$ is the electron energy at time $t$ [Glauber_1951, Di_Piazza_2010]. Equation (1) can then be used to calculate the relative electron energy loss $\zeta$ defined as
$\zeta\equiv\frac{\mathcal{E}_{0}-\mathcal{E}_{F}}{\mathcal{E}_{0}},$ (2)
where $\mathcal{E}_{0}$ is the initial energy of an electron and $\mathcal{E}_{F}$ its energy after interacting with the laser pulse. Equation 1 applies to a single electron but can be interpreted as the continuous, average energy loss per-electron for an ensemble of electrons sharing the same initial condition. This is verified in Section I of the Supplemental Material (SM) [supplemental] by comparing the results of Eq. (1) to a stochastic quantum mechanical approach (see also Ref. [niel2018quantum]).
[…]
The advantage of using a FF pulse in place of a SFG pulse with the same energy will be demonstrated by numerically integrating Eq. (1). An analytical treatment is also possible by making a few simplifying assumptions: (1) $\mathcal{E}\gg m\xi_{0}$, which guarantees that, for an electron initially counterpropagating with respect to the laser pulse, the transverse electron motion and pulse structure can be ignored; (2) The focal point of the laser pulse is located at $x=y=0$ and the electron moves along the negative $z$ axis at an ultrarelativistic velocity such that $z(t)\approx z(0)-t$; (3) The temporal profile of the pulse has a square shape equal to unity with a duration $2t_{\text{int}}$, which intersects the electron from $z=t_{\text{int}}/2$ to $z=-t_{\text{int}}/2$.
In the classical regime ($\chi\ll 1$), Eq. (1) can be approximated as $d\mathcal{E}/dt\approx-2\alpha m^{2}\chi^{2}/3$. When interacting with the FF pulse, the electron experiences a field with a near-constant amplitude $\xi_{0}$ such that $\chi_{\text{FF}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}$. The relative energy loss [Formanek:2021bpw] is then
$\zeta_{\text{C,FF}}=\frac{\kappa_{\text{C,FF}}}{1+\kappa_{\text{C,FF}}},\>\>\kappa_{\text{C,FF}}\approx 2\frac{U(\text{J})\mathcal{E}_{0}(\text{GeV})}{[\sigma_{0}(\mu\text{m})]^{2}}\,$, (3)
where $U=\pi E_{0}^{2}\sigma_{0}^{2}t_{\text{int}}/2$ is the pulse energy [Esarey:1993zz] and $\sigma_{0}$ its focal spot size. This result shows that the energy loss in the classical regime depends on the fluence of the laser pulse, i.e., it is $\propto U/\sigma_{0}^{2}$, but is independent of the interaction time $t_{\text{int}}$  [^1] Note that subleading oscillatory terms were ignored.. For a SFG pulse, the finite Rayleigh range $z_{R}=\omega_{0}\sigma_{0}^{2}/2$ can be taken into account by substituting $\xi_{0}\rightarrow\xi_{0}/\sqrt{1+z^{2}/z_{R}^{2}}=\xi_{0}/\sqrt{1+t^{2}/z_{R}^{2}}$ in the expression for $\chi$, such that $\chi_{\text{SFG}}(t)\approx 2\mathcal{E}(t)\omega_{0}\xi_{0}|\sin(2\omega_{0}t)|/m^{2}\sqrt{1+t^{2}/z_{R}^{2}}$. As a result, $\kappa_{\text{C,SFG}}=\kappa_{\text{C,FF}}\arctan(\rho)/\rho$, with $\rho\equiv t_{\text{int}}/2z_{R}$ [Formanek:2021bpw]. Thus, for a fixed laser energy, a FF pulse and a short SFG pulse with $\rho\ll 1$ induce the same energy loss in the classical regime.
In the quantum regime ($\chi\gtrsim 1$), the scalings change because the emission rate [Eq. (1)] has a weaker dependence on the field amplitude than in the classical regime. This suggests that, in the quantum regime, a long FF pulse can induce more energy loss than a short SFG pulse with the same energy. This reasoning can be verified analytically for FF pulses in the limit $\chi\gg 1$, where $d\mathcal{E}/dt\propto-\chi^{2/3}$ and
$\mathcal{E}_{0}^{1/3}-\mathcal{E}_{F}^{1/3}\propto U^{1/3}t_{\text{int}}^{2/3}.$ (4)
Thus, for a fixed pulse energy $U$, higher energy losses can be achieved by extending the interaction time.
Figure 1 confirms the above expectations by showing the relative energy loss as a function of the interaction time for electrons colliding head-on with either a FF pulse (blue solid lines) or a SFG pulse (dashed orange lines) with energies of (a) 1 J or (b) 10 J. The solid and dashed lines were obtained by numerically integrating Eq. (1), with $\chi$ equal to $\chi_{\text{FF}}$ or $\chi_{\text{SFG}}$, respectively. In both cases, the maximum value of the quantum nonlinearity parameter is $\chi_{0}=2\mathcal{E}_{0}\omega_{0}\xi_{0}/m^{2}\approx 0.1\,\xi_{0}\mathcal{E}_{0}(10\;\text{GeV})$. For $U=$ 1 J, the maximum energy loss achieved by a FF pulse is approximately $30\%$ higher than that of an SFG pulse regardless of the initial electron energy. For $U=$ 10 J and $\mathcal{E}_{0}=$ 0.5 GeV, the maximum energy loss achieved by a FF pulse is $12\%$ higher. In this case and for larger electron energies, the enhanced energy loss in the FF pulses is reduced because the electron loses almost all of its energy whether it interacts with a FF or SFG pulse (see also the inset of Fig. 2a discussed below).
Figure 1 also shows that the energy loss in SFG pulses is suppressed for either very short pulses ($\rho\ll 1$) or very long pulses ($\rho\gg 1$). Shorter duration SFG pulses have higher peak intensities, but, as discussed above, the emission rate has a weaker scaling with intensity than it does with duration in the quantum regime [niel2018quantum]. With long duration SFG pulses, the electron spends a large fraction of the interaction time outside of the confocal region where the intensity is lower. This contrasts with FF pulses where the electron spends the entire interaction near the focus where the intensity is high. As a result, the energy loss increases monotonically with the interaction time until it approaches the classical limit [green dotted lines evaluated using Eq. (3)]: for fixed pulse energy, $\chi_{0}\propto t_{\text{int}}^{-1/2}$. The energy loss in the FF and SFG pulses coincide at short interaction times because the two pulses have the same amplitude and the finite Rayleigh range of the SFG pulse does not play a role ($\rho\ll 1$).
The results presented in Fig. 1 were verified by evolving electrons in the full FF and SFG fields with a Monte Carlo radiation emission calculation that takes into account the stochastic nature of multiple photon emission (see Section I of SM [supplemental]). Notwithstanding, some remarks are in order to address the validity of the LCFA. The LCFA overestimates the amplitude of the exact QED emission spectrum for sufficiently low photon energies [Di_Piazza_2018, Di_Piazza_2019, Ilderton_2019_b]. The key parameter is [Di_Piazza_2018]
$\eta_{\text{LCFA}}=\frac{\mathcal{E}-\omega}{\omega}\frac{\chi}{\xi^{3}}\,$, (5)
where $\omega$ is the energy of the emitted photon and $\xi$ and $\chi$ are the classical and quantum nonlinearity parameters at the instant of emission. The condition $\eta_{\text{LCFA}}^{2/3}\ll 1$ ensures that corrections to the LCFA are small [Di_Piazza_2019]. Equation (1) is derived by integrating the emitted photon spectrum over the photon energy. For $\chi\ll 1$, the largest contribution to the integral comes from energies $\omega\lesssim\chi\mathcal{E}\ll\mathcal{E}$; for $\chi\sim 1$, the largest contribution comes from energies $\omega\lesssim\mathcal{E}$. As a result, the validity of the LCFA is determined by the low-energy photons and the condition $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi^{2}\ll 1$. The vertical line in Fig. 1 corresponds to the choice $\xi_{0}=5$ [Blackburn:2018sfn] so that $\eta_{\text{LCFA}}^{2/3}=1/\xi_{0}^{2}=0.04$. Note that the expression for the classical energy loss does not rely on the validity of the LCFA, which explains why in the FF case the curves tend to the classical result as the interaction time increases ($\chi_{0}\propto t_{\text{int}}^{-1/2}$ for fixed energy).
[…]
To further illustrate the advantage of FF pulses in the quantum regime of radiation reaction, Fig. 2 compares the maximum energy loss induced by FF pulses (with the longest $t_{\mathrm{int}}$ satisfying $\eta_{\text{LCFA}}^{2/3}\sim 1/\xi_{0}^{2}<0.04$) to that of SFG pulses (the maxima of the orange dashed curves in Fig. 1). For almost all initial electron energies $\mathcal{E}_{0}$, the FF pulses (solid lines) result in a larger loss of electron energy than the SFG pulses (dashed lines). However, as the pulse energy increases, the electron energy loss for both the FF and SFG pulses approaches unity (Fig. 2a), which limits the relative advantage of the FF pulses (Fig. 2a inset). In addition, both pulses achieve higher energy losses with smaller spot sizes (Fig. 2b) [see also Eq. (3) for the classical regime]. Thus, defocusing an SFG pulse to increase the interaction time would not improve its results.
Figure 3: Number of photons emitted per electron at $\chi>0.1$ from the on-axis collision of $\mathcal{E}_{0}=10$ GeV electrons with 1 J FF and SFG laser pulses. The inset shows the number of photons emitted as a function of energy and $\eta_{\text{LCFA}}^{2/3}$ for the FF pulse. The numbers in white indicate what percentage of photons in the given range were emitted at $\chi>0.1$.
II.1 Photon yield
The saturation of the electron energy loss at higher pulse energies suggests that an observable without a theoretical upper limit would better elucidate the benefits of FF pulses in these conditions. One such observable, that is also important for applications, is the total number of emitted photons. Before presenting numerical calculations of the photon number, it is instructive to examine its scaling behavior to explain the advantages of a FF pulse compared to a SFG pulse. Within the LFCA, the probability $P$ of single photon emission per unit time can be calculated using the interpolation formula [Baier:1998vh]
$\frac{dP}{dt}\approx\frac{1.45\alpha m^{2}\chi}{\mathcal{E}(1+7.2\chi+\chi^{2})^{1/6}}\,$, (6)
which is better than 2% accurate for all values of $\chi$. Multiplying the time-integrated probability by the number of electrons colliding with the laser pulse provides an estimate of the average number of emitted photons. Provided that the energy loss is accounted for as discussed below Eq. (1), this estimate is quantitatively accurate if higher-order processes, such as multiple photon emission within a formation length can be neglected.
In the classical regime ($\chi\ll 1$), Eq. (6) scales as $dP/dt\propto\chi/\mathcal{E}$; while in the quantum regime ($\chi\gg 1$), $dP/dt\propto\chi^{2/3}/\mathcal{E}$. These proportionalities indicate that the emission probability $P$ scales more favorably with the interaction time than the laser intensity in both regimes. Thus, for fixed collision parameters (i.e., pulse energy, electron energy, focusing), FF pulses have an advantage over SFG pulses regardless of $\chi$.
To demonstrate this, we first consider the most favorable conditions for a SFG pulse, which is the classical regime where $\chi\ll 1$. Maximizing the emission probability for a SFG pulse requires matching the interaction time $t_{\text{int}}$ to the Rayleigh range $z_{R}$, so that electrons spend much of the interaction near the focus where the intensity is high. The optimum is $t_{\text{int}}\approx 5z_{R}-7z_{R}$, which yields $P_{\text{C,SFG}}\propto U^{1/2}$ (see Section II of SM [supplemental]). In a FF pulse, the focus moves with the electrons, so the entire interaction time is spent near the focus. As a result, the interaction time is independent of the FF Rayleigh range $z_{R,\text{FF}}\equiv\omega_{0}\sigma_{0}^{2}$. In this case, $P_{\text{C,FF}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}U^{1/2}$, such that the average number of emitted photons increases with the interaction time. The ratio $P_{\text{C,\text{FF}}}/P_{\text{C,SFG}}\propto(t_{\text{int}}/z_{R,\text{FF}})^{1/2}$ reveals that, for fixed laser pulse energy, a FF pulse will have an advantage over a SFG pulse when the interaction time of the FF is longer than its Rayleigh range. In the deep quantum regime ($\chi\gg 1$) the scaling is even more favorable, starting as $P_{\text{Q,FF}}\propto U^{1/3}t_{\text{int}}^{2/3}$ before asymptoting to the classical scaling as the electron loses energy.
Quantum enhancement.
Figure 3 demonstrates that a FF pulse can yield more photons emitted in the quantum regime than SFG pulses with the same energy. In the on-axis collision of 10 GeV electrons with a 1 J, $\xi_{0}=6$ FF pulse (as considered in Fig. 1a), 10.0 photons were emitted at $\chi>0.1$ per electron in the energy range 100 MeV – 2 GeV compared to 5.2 photons per electron with a 1 J, $\xi_{0}=13$ SFG pulse (maximum energy loss from Fig. 1a) and compared to 6.3 photons per electron in a 1 J, $\xi_{0}=9$ SFG pulse. The amplitude, spot size, and interaction time ($t_{\text{int}}=128$ fs) of the $\xi_{0}=9$ SFG pulse were optimized to achieve the highest photon yield with $\chi>0.1$ in the 100 MeV – 2 GeV range. The electrons lost 84%, 64%, and 59% of their initial energy in the FF, $\xi_{0}=13$ SFG, and $\xi_{0}=9$ SFG pulses, respectively. The electron energy distributions had initial standard deviation widths of 1%, which were broadened by quantum photon emission to 39%, 50%, and 38% with maximum quantum parameters $\chi_{0}\approx 0.59$, $1.23$, and $0.83$ [Neitz_2013, Neitz_2014, Vranic:2015sft]. The photon spectra were generated using the Monte Carlo code described in Section I of SM [supplemental]. Photons emitted with energies below 100 MeV in the FF collision are not reported because they do not meet the requirement for satisfying the LCFA approximation discussed below Eq.(5).
Photon yield increase.
[…]
Figure 4 shows that the FF pulse colliding with 10 GeV electrons results in more photons per electron than the SFG pulse by a factor of 7 in the 1 – 2 MeV range, a factor of 5 in the 2 – 10 MeV, and a factor of 4 in 10 – 20 MeV range. Based on these results, the collision of a 100 pC beam of $\mathcal{E}_{0}$ = 10 GeV electrons with a FF pulse could produce $5\times 10^{10}$ photons in the 1 – 20 MeV range, compared to only $9\times 10^{9}$ photons with a SFG pulse.
Despite its much lower amplitude ($\xi_{0}=6$), the FF pulse produces a substantial number of high-energy photons ($\omega_{0}>20$ MeV) comparable to that produced by the higher amplitude ($\xi_{0}=16$) SFG pulse. However, most of the photons are in the MeV range. For 10 GeV electrons, there are two-times more photons in the 1–20 MeV range (accounting for 5% of $\mathcal{E}_{0}$ on average) than above 20 MeV (accounting for 92% of $\mathcal{E}_{0}$). The maximum quantum parameter in the FF collision is $\chi_{0}\approx 0.6$ and many photons are emitted in the quantum regime of radiation reaction.
For a FF pulse colliding with 1 GeV electrons, there are eight-times more photons in the 1–20 MeV range than above 20 MeV. The 1–20 MeV photons account for 36% of $\mathcal{E}_{0}$ on average, while the photons above 20 MeV account for 34%. The higher photon yield produced by the FF is due to a larger number of photons below 20 MeV. As $\chi_{0}\approx 0.06$, the classical scaling applies, and the 3 times increase in the number of 1 – 20 MeV photons is comparable to the $\sim$2.7 times overall increase predicted by the classical scaling, $dP/dt\propto\chi/\mathcal{E}$.
Figure 4: Spectra of emitted photons from the head-on collision of $\mathcal{E}_{0}=10$ GeV and 1 GeV electrons with $U=10$ J laser pulses. The inset shows the number of photons emitted as a function of energy and $\eta_{\text{LCFA}}^{2/3}$ for the FF pulse collisions. The number of photons emitted per electron in each shaded interval is displayed in the bottom left. The FF pulse (solid lines) results in approximately five-times and three-times more photons in the 1–20 MeV range than the SFG pulse (dashed lines) for 10 and 1 GeV electrons, respectively.
III Simulation validation
Because the advantage of the FF pulse occurs in the low-energy part of the spectrum, it is important to verify that the breakdown of the LCFA at low photon energies does not invalidate the results [see Eq. (5) and Refs. [Di_Piazza_2018, Di_Piazza_2019, Ilderton_2019_b]]. To assess possible deviations from the LCFA in this part of the spectrum, simulations were performed with the particle-in-cell (PIC) code Smilei [DEROUILLAT2018351] including the SFQED-tookit [Montefiori:2023mkc], which implements corrections beyond the LCFA [Di_Piazza_2019]. These simulations are in excellent agreement with the Monte Carlo results presented in Figs. 3 and 4 (see Section III of SM [supplemental]). As additional verification, the insets of Figs. 3 and 4 show that almost all photons in the studied ranges are emitted with $\eta_{\text{LCFA}}^{2/3}<0.1$.
Finally, the Monte Carlo code used to generate the photon spectra in Figs. 3 and 4 does not account for the decay of the emitted photons into electron-positron pairs, which would also radiate [mercuri2021impact]. As demonstrated in Section IV of the SM [supplemental], pair production can be safely neglected for both the SFG and FF pulses considered in Figs. 3 and 4. At higher field strengths ($\xi_{0}\gtrsim 40$), pair production in a SFG pulse would deplete the most energetic photons and slightly increase the low-energy photon yield, but the overall energy loss and photon yield would be suppressed (see Section III of the SM [supplemental]).
IV Conclusions
In conclusion, we have shown that a flying focus (FF) pulse can significantly enhance the signatures of quantum radiation reaction when compared to a stationary focus Gaussian (SFG) pulse with the same energy. First, an electron can radiate significantly more energy in a FF pulse than in a SFG pulse. Second, a FF pulse enhances the yield of emitted photons, especially in the low-energy part of the spectrum. Moreover, the number of photons emitted in the quantum regime of the interaction (with $\chi>0.1$) can be larger when using a FF in place of a SFG pulse. These improvements are a direct result of the energy loss in the quantum regime and the photon emission probability in both the quantum and classical regimes scaling more favorably with the interaction time than the field intensity. Thus, any Compton scattering configuration can use a FF pulse in place of a SFG pulse to exploit this advantage. Furthermore, the use of FF pulses lowers the required power and intensity, bypasses the need for larger compression gratings, and provides a more controllable environment with simpler diagnostics of the interaction field strength [Popruzhenko:2023tqx], while simultaneously enhancing the radiation signatures.
Acknowledgments. The authors thank Samuele Montefiori and Matteo Tamburini for kindly sharing a development version of their SFQED-toolkit module [Montefiori:2023mkc] for SMILEI and providing support. We also thank Elias Gerstmayr for discussions regarding laboratory sources of gamma rays. This project has received funding from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101105246-STEFF. This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90254). This material is based upon work supported by the U.S. Department of Energy [National Nuclear Security Administration] University of Rochester “National Inertial Confinement Fusion Program” under Award Number DE-NA0004144 and U.S. Department of Energy, Office of Science, under Award Number DE-SC0021057.
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
See pages 1 of supplementalSee pages 2 of supplementalSee pages 3 of supplementalSee pages 4 of supplementalSee pages 5 of supplementalSee pages 6 of supplemental